SM.51 The Candy Shack is in the process of finding suppliers for their candy stands. They have determined to single-source their supply in order to focus on the buyer/vendor relationship. The Candy Shack has narrowed their supplier list to two candidates: Gem State, a local company in Pocatello, and Metro Candy, an online wholesaler. The table below contains the Candy Shack's weighted criteria for making this decision along with the raw scores for each of the two supplier options.

Criteria Weight   Gem State   Metro Candy  
Price 0.20 9 8
Quality 0.25 9 8
Lead Time 0.25 7 6
Terms 0.15 3 7
Reputation 0.15 8 8


What is Gem State's weighted score for PRICE? (Display your answer to two decimal places.)
   

What is Metro's weighted score for PRICE? (Display your answer to two decimal places.)
   

What is Gem State's weighted score for QUALITY? (Display your answer to two decimal places.)

   

What is Metro's weighted score for REPUTATION? (Display your answer to two decimal places.)

   

What is Gem State's TOTAL weighted score? Display your answer to two decimal places.)

   

What is Metro's TOTAL weighted score? Display your answer to two decimal places.)
Answer #1

Weighted Score of Criteria = Weight of the Criteria * Rating of Respective Criteria for given optios

Gem State's weighted score for PRICE = 0.20*9 = 1.80

Metro's weighted score for PRICE = 0.20*8 = 1.60

Gem State's weighted score for QUALITY = 0.25*9 = 2.25

Metro's weighted score for REPUTATION = 0.15*8 = 1.20

Gem State's TOTAL weighted score = 0.20*9 + 0.25*9 + 0.25*7 + 0.15*3 + 0.15*8 = 7.45

Metro's TOTAL weighted score = 0.20*8 + 0.25*8 + 0.25*6 + 0.15*7 + 0.15*8 = 7.35